Manuel Anselmo is ten times older than his son, Andrew. In 20 years, Manuel's age will be 4 more than twice Andrew's age. How old was he when Andrew was born?
Found 2 solutions by mananth, amoresroy:
Answer by mananth(16946)

Andrew = x
father = 10x
..
after 20 years
--------------
son x+20
father 10x+20
..
10x+20 = 2(x+20)+4
10x+20=2x+40+4
10x-2x=40-20+4
8x=24
/8
x= 3 andrew's age now
father's age = 3x = 30
..
3 years ago when andrew was born father was 30-3= 27 years.
...
m.ananth@hotmail.ca

Answer by amoresroy(361)

Manuel Anselmo is ten times older than his son, Andrew. In 20 years, Manuel's age will be 4 more than twice Andrew's age. How old was he when Andrew was born?
Let a = the age of Andrew now
20a = the age of Manuel now
20a-a = the age of Manuel when Andrew was born
20a + 20 = 2(a+20) + 4
20a + 20 = 2a+40 + 4
18a = 24
a = 4/3
20a = 80/3
20a-a = (80-4)/3
= 76/3 = 25 and 1/3
Manuel was 25 and 1/3 years old when Andrew was born